OpenStudy (anonymous):

Determine if the following Serie converges or not \[\sum_{k\in \mathbb{N}}2^{2-k}\] is the solution : yes its convergence and convergt against 8 ?

OpenStudy (anonymous):

\[\sum_{k \in \mathbb{N}}2^{2-k}=4\sum_{k \in \mathbb{N}}\frac{ 1 }{ 2^k }=4\cdot 2 = 8\]

OpenStudy (anonymous):

that means the solution right sure? thank you very much pgpilot

OpenStudy (anonymous):

it converge right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you're welcome